Laci:

John has 54 white marbles and 9 yellow marbles. Ben has 16 white marbles and 96 yellow marbles. Noel has 12 white marbles and 72 yellow marbles. The boys who have the same ratio of white marbles to yellow marbles are .

Gdeinward:

Any ideas?

Laci:

There is 4 options A.) John and Noel B.) John and Ben C) Noel and Ben D.) All I thought it was C, but I have no idea

Gdeinward:

are you familiar with GCF, you know to simplify fractions?

Laci:

I know how to simplify fractions, but i"m not that great at GCF, but I can kinda do it

Laci:

Nevermind I got A

Gdeinward:

Well you simply must do the same with ratios for example: 54:9 = 54/9 and 9/ 9 = 6/1= 6

Laci:

That"s punctually what I just did I re-read it and thought about what you said about fractions so I made them into factions and divided all of the and got 6 for both John and Noel

Gdeinward:

That is incorrect. 1/6 does not equal six

Gdeinward:

Its Ben and John

Laci:

I thought you would divided each persons by their colors like 71 divided by 12 and 54 by 9

Gdeinward:

No, look at how I completed the example. you divide each number by a number that factor it into a whole number

Laci:

I don't see anything,but i'm starting to get it

Gdeinward:

16:96 = 16/96 = 16 /4 = 4. 96/4 = 24. so 4/24. 4/24 = 1/6.

Laci:

Oh Ok Thank You So Much

Gdeinward:

Yw. Once you get it down, it is quite simple. I wish you the best of luck!

Laci:

Thank You So Much,You Have Been The Best HelpπŸ˜€πŸ˜πŸ‘